OpenStudy (anonymous):

The shoerter leg of a 30-60-90 right triangle meadsures 3 in whats the length of the hypothenus.

OpenStudy (anonymous):

\[a) 3\in b) 3\sqrt{3} c) 3\sqrt{3} d) 9\]

OpenStudy (anonymous):

It's been a long time since geometry and I don't like it much but I think you can set it up as cos(30) = 3/h

OpenStudy (anonymous):

uh lol

OpenStudy (anonymous):

In triangle ABC, AB = 13 cm, BC = 25 cm, and AC = 17 cm. What is the smallest angle of triangle ABC? C B The smallest angle of triangle ABC can not be determined with the given information. A

OpenStudy (anonymous):

angle c b or a. is what that meant

OpenStudy (anonymous):

The smallest angle is angle C. A picture really helps with this one.

OpenStudy (anonymous):

PQS is only similar to RQP. RQP is only similar to RPS. PQS is similar to both RQP and RPS. None of the triangles are similar

OpenStudy (anonymous):

a b c or d

OpenStudy (anonymous):

how do you kno this stuff?

OpenStudy (anonymous):

I can't see that one. Similar triangles mean the angles are the same, but the sides may have different lengths (that are proportional).

OpenStudy (anonymous):

you can enlarge the image by presein it

OpenStudy (anonymous):

I know, but I can't view it at work (it won't let me)

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

thanks anyhow

OpenStudy (anonymous):

thanks anyhow

OpenStudy (anonymous):

Sorry I'm not more help, don't do much geometry.